OpenStudy (anonymous):

find all real solutions of the equation 3x+12=24

OpenStudy (anonymous):

3x+12=24 -12 -12 3x=12 x=4

OpenStudy (anonymous):

really?? thats it???

OpenStudy (anonymous):

yeah